OpenStudy (anonymous):

what is the radius? x^2+y^2-2x+6y+9=0

OpenStudy (anonymous):

\[\sqrt{9}=?\]

OpenStudy (usukidoll):

maybe it helps to rearrange this equation first and then have it in the equation of the circle format.

OpenStudy (freckles):

completing the square needs to be done for the x and y terms separately of course

OpenStudy (usukidoll):

OH right forgot that @freckles

OpenStudy (anonymous):

oh yes.

OpenStudy (anonymous):

the answer is 1 thank you all though :)

OpenStudy (usukidoll):

how did you get 1?

OpenStudy (anonymous):

graphing calculator perhaps

OpenStudy (usukidoll):

or desmos/wolfram alpha XDDDDDDDDDDDDD

OpenStudy (anonymous):

\[(x^2-2x)+(y^2+6y)=-9\]

OpenStudy (anonymous):

i didnt but i got answer wrong and the test said the correct anwer was 1 lol

OpenStudy (anonymous):

thats why we never get mcq for test

OpenStudy (usukidoll):

<_< yeah that's why I asked... you need to understand the concepts of this so you can be able to solve these equations

OpenStudy (usukidoll):

mcq? @saseal

OpenStudy (anonymous):

multiple choice questions

OpenStudy (usukidoll):

oh XD. Come to think of it, it's been years since I had one

OpenStudy (anonymous):

@grecar98 take that equation i gave ya and complete the square

OpenStudy (anonymous):

time for me to learn bitumen road making...afk!

OpenStudy (irishboy123):

as the OP has found the answer: \(x^2+y^2-2x+6y+9=0\) rearrange \(x^2 - 2x +y^2+6y+9=0\) complete squares \([(x-1)^2 -1] +[(y +3)^2 - 9] +9=0\)